ENHZ:

Tickets for a school play cost $4 for adults and $2 for students. At the end of the play, the school sold a total of 105 tickets and collected $360. Use the system of equations you wrote to find the number of adult tickets sold and the number of student tickets sold.

ENHZ:

x + y = 105 4x+2y=360

ENHZ:

@AZ you are the only smart person I know that is smart

AZ:

do you want to solve the system of equations by substitution, elimination or by graphing?

ENHZ:

Umm. I need a (x, y) answer

AZ:

yes, so we have to solve for x and for y and then we write the two numbers like (x, y)

AZ:

Do you know what the substitution or elimination method is?

ENHZ:

Not really. .-.

AZ:

okay so we have 2 equations let's solve the first equation for x x + y = 105 if we subtract y on both sides, what do you get?

ENHZ:

would it be -105x?

AZ:

no x + y = 105 we SUBTRACT y on both sides so x + y - y = 105 - y what is y - y =

ENHZ:

0

AZ:

it's like saying I have 1 apple and I give away the apple, how many apples do I have

AZ:

Good!

AZ:

so x = 105 - y does that make sense?

ENHZ:

yeah!

AZ:

so now we substitute this equation into the other one so we plug it back in so we have x = 105 - y and 4x+2y=360 since x equals 105-y then that means in the second equation where we have 'x', we can replace it with 105-y because they're the same thing so we get 4x + 2y = 360 4(105-y) + 2y = 360 can you distribute the 4 into the parenthesis?

ENHZ:

im not sure i understand. do i do 4*105?

AZ:

yeah that's part of it so basically when we distribute, it's like \( a(b-c) = ab - ac\) so \( 4(105 - y)\) is going to be 4*105 - 4*y does that make sense?

ENHZ:

I think so. Would it be 420-4y?

AZ:

perfect! so now we have 420 - 4y + 2y = 360 what is -4y + 2y =

ENHZ:

-6y?

ENHZ:

oh my bad, its -2y

AZ:

aha- there you go!

ENHZ:

so, 420-2y=360

AZ:

so we have 420 - 2y = 360 so now what do you get if you subtract 420 on both sides

ENHZ:

-2y=(-60)

ENHZ:

x=30?

zunairah:

bruh this was on my uesterdays hw

AZ:

be careful, you switched from y to x but yes y = 30

ENHZ:

oh ok

AZ:

so good job!! and now that we know what y is, we can solve for x we know that x + y = 105 so if y = 30 then what is x?

ENHZ:

75?

Ashtag:

you don't get this kind of collaboration in class that's for sure

ENHZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Ashtag you don't get this kind of collaboration in class that's for sure \(\color{#0cbb34}{\text{End of Quote}}\) shush

ENHZ:

so (75, 30)

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ENHZ 75? \(\color{#0cbb34}{\text{End of Quote}}\) good job so x = 75 and y = 30 so write it as (x, y) now

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ENHZ so (75, 30) \(\color{#0cbb34}{\text{End of Quote}}\) splendid job :))

ENHZ:

yay ty @AZ

Ashtag:

poggies

ENHZ:

took me long enough

AZ:

No problem!

ENHZ:

only took 45 replies

Ashtag:

now its 45